OpenStudy (anonymous):

Write the equation of the line that passes through (2, 3) and is parallel to the line 12x – 5y = 2.

OpenStudy (anonymous):

Parallel lines have same slope .... Slope of 12x-5y-2=0 is -(coefficient of x)/(coefficient of y) ie -12/-5 ie 12/5 Thus equation of line: y-3=(12/5)(x-2)

OpenStudy (anonymous):

SO IS IT A